A cell divides into two identical copies every 4 minutes. How many cells will exist after 5 hours?

Answer by stanbon(75887)

A cell divides into two identical copies every 4 minutes. How many cells will exist after 5 hours?
----------------
A(t) = Ao*(2)^(t/4 min)
-----------------------
Note: 5 hours = 5*60 = 300 minutes
-----------------------
A(300) = 1*2^(300/4)
A(300) = 2^75
A(300) = 3.777893186..x 10^22 cells
=========================
Cheers,
Stan H.